In a quadrilateral OABC, D is the midpoint of BC and E is the point on AD such

that AE : ED = 2 : 1. Given that OA = a, OB = b and OC = c, express OD and

OE in terms of a, b and c.


1
Expert's answer
2021-02-03T03:39:49-0500

Solution.

By the rule of adding vectors we get

"\vec{OA}+\vec{AB}=\vec{OB},"

from here "\vec{AB}=\vec{OB}-\vec{OA}=b-a."

By the rule of adding vectors we get

"\vec{OB}+\vec{BC}=\vec{OC},"

from here "\vec{BC}=\vec{OC}-\vec{OB}=c-b."

"D" is the midpoint of "BC", then "\vec{BD}=\frac{c-b}{2}."

By the rule of adding vectors we get

"\vec{OD}=\vec{OB}+\vec{BD}=b+\frac{c-b}{2}=\frac{b+c}{2}."

By the rule of adding vectors we get

"\vec{AD}=\vec{AB}+\vec{BD}=b-a+\frac{c-b}{2}=\frac{b+c-2a}{2}."

"E" is the point on "AD" such that "AE:ED=2:1," then "\vec{AE}=\frac{2}{3}\vec{AD}=\frac{2}{3}\frac{b+c-2a}{2}=\frac{b+c-2a}{3}."

By the rule of adding vectors we get

"\vec{OE}=\vec{OA}+\vec{AE}=a+\frac{b+c-2a}{3}=\frac{a+b+c}{3}."

Answer.     

"\vec{OD}=\frac{b+c}{2}"

 "\vec{OE}=\frac{a+b+c}{3}"